You have started working as a cost accountant for a firm that has only been in business for one month. The firm is able to buy a new type of biodegradable plastic at a fixed price of $100 per roll. The plastic is then cut and sealed to make garbage bags. Fixed factory overhead is estimated to be $125,000 per month. During this past month, 8,000 cartons of garbage bags were produced, which represents 80% of the activity volume. You are given the following information:

Rolls of plastic used 40
Variable overhead incurred $61,000
Overhead efficiency variance $5,000 U
Standard costs per carton of garbage bags:
Labour hours 2
Wage rate $8 per hour
Total overhead $20
Rolls of plastic 0.004 rolls

a. Applied overhead per direct labour hour.

b. Standard direct labour hours allowed for unit produced

c. Activity volume

d. Predetermine fixed overhead rate

e. Fixed overhead applied

f. Variable overhead spending variance

g. Actual number of direct labour hours incurred

h. Labour efficiency variance

i. Materials quantity variance

j. Fixed overhead budget variance

k. Fixed overhead volume variance
Answer #1
a. Applied overhead per direct labour hour
$ 20/2 hrs.= 10 per carton
b. Standard direct labour hours allowed for unit produced
2 hrs.*8000 cartons = 16000 hrs.
c. Activity volume
80%(Given)
d. Predetermined fixed overhead rate
Estimated FOH p.m.= $ 125000
100% activity volume= 8000/80%*100%=
10000
Cartons
So, Predetermined FOH /Carton=
$ 125000/10000 Cartons =
12.5
per carton
e. Fixed overhead applied
$ 12.5*8000 cartons=
100000
f. Variable overhead spending variance
VOHSV
Total OH/carton(Std.).given as ------------ $ 20
As per d. above, Std. fixed OH rate/carton=$ 12.5
So, Std. VOH rate/carton = (20-12.5)= $ 7.5 /carton
VOHSV= (Std.VOH rate-Actual VOH rate)/Carton*Actual cartons
(7.5-(61000/8000))*8000=
-1000
UF
g. Actual number of direct labour hours incurred
VOHEff. V=(Std. hrs.for actual prodn.-Actual Hrs.)*Std.Rate
-5000=((2*8000)-Actual hrs.)*7.5
Actual hrs.= 16666.7 hrs.
h. Labour efficiency variance
LEV=(Std. hrs.for actual prodn.-Actual Hrs.)*Std. wage Rate
((2*8000)-16667)*8=
-5336
UF
i. Materials quantity variance
MQV=(Std. qty. for actual prodn.-Actual qty.)*Std.price/qty.
ie.((0.004*8000)-40)*100=
-800
UF
j. Fixed overhead budget variance
FOH BV=Total FOH budgeted-Actual FOH incurred
125000-125000=0
k. Fixed overhead volume variance
FOHVV=Absorbed/applied FOH-Budgeted FOH
(8000*12.5)-125000=
-25000
UF